A projects payback period (PB) indicates the number of years required for a project to recover its initial investment using its operating cash flows. As the theoretical soundness of the conventional (undiscounted) PB technique was criticized, the model was modified to incorporate the time value of money-adjusted operating cash flows to create the discounted payback method. While both payback models continue to reflect faulty ranking criteria, they do provide important (useful) information regarding a projects liquidity and riskiness. In general, the shorter the payback, other things constant, the greater the projects liquidity. Suppose you are evaluating a project with the expected future cash inflows shown in the following table. Your boss has asked you to calculate the projects net present value (NPV). You dont know the projects initial cost, but you do know the projects regular, or conventional, payback period is 2.50 years. If the projects mWACC~ is 10%, the projects NPV (rounded to the nearest dollar) is: Year Cash Flo Year $300,000 Year 2 $475,000 Year 3 $425,000 Year 4 $450,000 O $334,899 O $304,454 O $319,677 $365,345 Which of the following statements indicate a disadvantage of using the regular payback period (not the discounted payback period) for capital budgeting decisions? Check all that apply. The payback period does not take the projects entire life into account. The payback period does not take the time value of money into account. The payback period is calculated using net income instead of cash flows.Along with the questions, is the first one shorter or longer?
Answer #1

Shorter the payback, greater is the liquidity.

So, the first answer "Shorter" is correct.

Question 1

Conventional payback period refers to simple payback period.

In this case, simple payback period is 2.5 years. This means the project has taken 2.5 years to recover it's initial investment.

Cash generated by the project in 2.5 years = C1 + C2 + C3 / 2 = 300,000 + 475,000 + 425,000 / 2 = 987,500

Please note that i have assumed only half of cash flow in year 3 (represented by C3) will be realized by the middle of year (i.e. year 2.5).

Hence, the initial investment in the project = C0 = -987,500

Discount rate = WACC = 10%.

Please see the table below for NPV calculation. The linkage column explains how each row has been calculated.

Year, N

0

1

2

3

4

Cash flows

Ci

(987,500)

300,000

475,000

425,000

450,000

Discount rate

WACC

10%

PV factor

(1 + WACC)(-N)

1.0000

0.9091

0.8264

0.7513

0.6830

PV of Cash flows

C x PV factor

(987,500)

272,727

392,562

319,309

307,356

NPV

Sum of all PV

304,454

Hence, the correct answer is second option representing $ 304,454

Question 2

The regular payback period doesn't take into account the cash flows occurring over the entire life of the project. It doesn't take into account the time value of money.

Hence, first and second options both are correct and should be ticked.

The third option is wrong. Payback period takes into account the cash flows and not accounting income.